OpenStudy (sydthekid913):

The solution to an absolute value inequality is 2003-07-04-00-00_files/i0230000.jpg. If the form of the absolute value inequality is 2003-07-04-00-00_files/i0230001.jpg, what is a? A. 7 B. 1 C. 14 D. 2

OpenStudy (sydthekid913):

please help

OpenStudy (sydthekid913):

Sorry it glitched here is the first equations \[-6\le x \le8\]

OpenStudy (sydthekid913):

and here is the second \[\left| x-a \right|\le b\]

OpenStudy (anonymous):

|x-a|\(\le \) b u can write it like this -b \(\le \) (x-a)\(\le \) b a-b \(\le \) x \(\le \) a+b -6 \(\le \) x \(\le \)8 soooo leavsu with two eq :- a-b=-6 a+b=8 2a=2 -----> a=1 :D

OpenStudy (anonymous):

have a nice day !

OpenStudy (sydthekid913):

cool thank u soooooooooo much

OpenStudy (anonymous):

np :D

OpenStudy (sydthekid913):

oooooooooohhhhhhhhh thanks for the medal!!!

OpenStudy (anonymous):

;-)